OpenStudy (anonymous):

Compute the derivatives 1. y = e^x ln x , (e rasied to the xlnx )

OpenStudy (lgbasallote):

\[\LARGE y = e^{x\ln x}\] first step ln both sides \[\LARGE \ln y = \ln e^{x\ln x}\] \[\LARGE \ln y = x\ln x\] do you see what i did?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

next step is implicit differentiation

OpenStudy (anonymous):

Ok what does that mean

OpenStudy (lgbasallote):

you take the derivative of the left side...and you take the derivative of the right side

OpenStudy (anonymous):

so I would be finding the derivative of e^x and lnx right?

OpenStudy (anonymous):

Careful....

OpenStudy (anonymous):

careful about what?

OpenStudy (anonymous):

Did you got what lgba told you???

OpenStudy (anonymous):

\(\LARGE \ln y = x\ln x\) Till here you got or not??

OpenStudy (anonymous):

yes I got it.

OpenStudy (anonymous):

Now take the derivative both the sides.. Do you know what is the derivative of logy???

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

\[\frac{d}{dy}Logy = \frac{1}{y}\] getting??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now take the derivative both the sides but implicitly.. Meaning by taking derivative on LHS it will become: \[\frac{1}{y}.dy\] Getting?? Now take the derivative on RHS.. Do you know about product rule in derivative??

OpenStudy (anonymous):

thats multiplying right?

OpenStudy (anonymous):

yes you have to use product rule there...

OpenStudy (anonymous):

\[\frac{1}{y}dy = (x.\frac{1}{x} + lnx).dx\] Getting till here??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now multiply both the sides by y what do you get?/ Show me..

OpenStudy (anonymous):

Multiply by y both the sides and tell me what you get??

myininaya (myininaya):

Or... \[y=e^{f(x)} => y'=f'(x)e^{f(x)}\]

OpenStudy (anonymous):

what happened last time? http://openstudy.com/users/hooverst#/updates/4fff7a07e4b09082c070a70c

OpenStudy (anonymous):

O lol dpalnc I had to leave and go pick up my mom from work

OpenStudy (anonymous):

That's what I am thinking @myininaya mam. But lgbasallote started it from other side so I was just explaining his method..

OpenStudy (anonymous):

I'm not sure if I'm solving it righ but would it be y=e^xlnx=y'(x)e^xlnx

OpenStudy (anonymous):

No its not..

OpenStudy (anonymous):

The correct form of this will be: \[\large y = e^{xlnx} \rightarrow y' = .e^{xlnx}.\frac{d}{dx}(x.lnx)\]

OpenStudy (anonymous):

Just apply the product rule to the last term...

OpenStudy (anonymous):

is the anserw e^xlnx=y=x^x

OpenStudy (anonymous):

Rewrite it in proper form..

OpenStudy (anonymous):

I have provided you the solution above so just apply the product rule there in the last term..

OpenStudy (anonymous):

ok so its y=x^x=e^xlnx ?

OpenStudy (anonymous):

No its not..

OpenStudy (anonymous):

See, tell me first: \[\large \frac{d}{dx}(x.lnx) = ??\] Apply product rule here and then show me..

OpenStudy (anonymous):

See, here in my 8th reply I have solved it you can see there.. My 8th reply..

OpenStudy (anonymous):

is it 2lnx/x* x^lnx-1

OpenStudy (anonymous):

How you got 2??

OpenStudy (anonymous):

Product rule says that: \[\large (uv) = u.v' + v.u'\] Using this you will get: \[\large \frac{d}{dx}(x.lnx) = x. \frac{1}{x} + lnx \rightarrow 1 + lnx\] So, \[\large \frac{d}{dx}(x.lnx) = 1 + lnx\] Now substitute this value there and then tell me what you got?

OpenStudy (anonymous):

You have to multiply one term with it just one term..

OpenStudy (anonymous):

\[y=e^xlnx \rightarrow y' =e^xlnx \times 1+lnx \] ?

OpenStudy (anonymous):

You are not getting yet.. See I have written as: \[\large y' = e^{xlnx} (\frac{d}{dx}(x. lnx))\] Now just substitute the value of \(\frac{d}{dx}(x.lnx)\) here and tell me what you get..